A is the point of intersection of the lines 3x − y = 7 and x + 4y + 2 = 0. Find
(a) in normal form, the equation of the line which passes through A and is parallel
to the line 2x + 3y − 40 = 0,
(b) the perpendicular distance of A from the line 2x + 3y − 40 = 0
1
Expert's answer
2019-11-06T10:41:35-0500

Find the point of intersection A by solving together two equations:"\left\{\begin{matrix}
3x-y=7
\\ 
x+4y=-2
\end{matrix}\right." . As a result obtain: "A=(2,-1)" .

Part (a):

  1. Find the equation of of the line which passes through A and is parallel to the line "2x + 3y \u2212 40 = 0". This two lines have common normal vector "b=(2,3)". The equation of the line passes through the point "(2,-1)" and has normal vector (2,3) is: "2(x-2) + 3(y+1) = 0 \Rightarrow 2x+3y-1=0".
  2. To transform this equation into the normal form one should divide both sides by the following expression: "\sqrt{2^2 + 3^2}". Thus obtain: "\dfrac{1}{\sqrt{13}}(2x+3y-1)=0 \Rightarrow \dfrac{2}{\sqrt{13}}x+\dfrac{3}{\sqrt{13}}y-\dfrac{1}{\sqrt{13}}=0".

Part (b):

  1. The distance between the point "(x_0,y_0)" and a line is given by the formula: "d=\dfrac{|ax_0+by_0+c|}{\sqrt{a^2 + b^2}}" , where "a,b,c" - coefficients in the line equation.
  2. With given data: "d=\dfrac{|2\cdot 2+3\cdot(-1)-40|}{\sqrt{13}} = \dfrac{39}{\sqrt{13}}".